OpenStudy (anonymous):

Is the set of multiples of 6 closed under subtraction? Explain why or provide a counterexample if not. A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12. @Haseeb96 @Punkchick @LazyBoy @satellite73 please help fan and medal.

OpenStudy (anonymous):

I think its A but I would double check

OpenStudy (anonymous):

k

OpenStudy (cwrw238):

no - not A because for example 36-12 = 24

OpenStudy (anonymous):

@cwrw238 is it b?

OpenStudy (anonymous):

@Lissa1199 do you think I'm right?

OpenStudy (anonymous):

I don't know I will look it up

OpenStudy (anonymous):

yes its b

OpenStudy (cwrw238):

to be honest I'm not sure whether 6 is classed as a multiple of 6 or not. If it is then the answer is c If not then the answer is b

OpenStudy (cwrw238):

i have forgotten a lot of this stuff can you for example include 12-12 as a subtraction ??

OpenStudy (haseeb96):

look here @Wolfie_GirlRocks2002 here Question is saying that the multiple of 6 is closed under subtraction ? look here i take two multiple number of 6 a) 12 b) 18 when we subtract 18-12=6 so 6 is also the multiple of 6 if a) 18 and b) 6 18-6=12 12 is also the multiple of 6 so 6 is closed under subtraction So the correct answer is B

OpenStudy (haseeb96):

we cant do that because in multiple of any number we never take 0

OpenStudy (haseeb96):

6 x 0 =0 it is also the multiple of 6

OpenStudy (haseeb96):

@cwrw238

OpenStudy (cwrw238):

oh ok

OpenStudy (haseeb96):

thanks dear